OpenStudy (anonymous):

Which of the following best describes the equation below? y = 4x + 10 A. linear B. both linear and nonlinear C. nonlinear D. neither linear nor nonlinear

OpenStudy (anonymous):

linear equation has no exponents

OpenStudy (anonymous):

Kelly works at her own nail salon. Today, she did 10 manicures and received $45.00 overall in tips. If x represents the cost of each manicure, which of the following equations can be used to find the total amount of money she earned, including tips, for her nail salon today? A. y = 10(x + 45) B. y = 45x + 10 C. y = 45(x + 10) D. y = 10x + 45

OpenStudy (anonymous):

let each manicure gives x dollars how much money she gets?

OpenStudy (anonymous):

I think its D... am I right?

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

x for 1 manicure how much for 10 manicures?

OpenStudy (anonymous):

@telltoamit

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The vertices of a trapezoid are shown below. P(0, 0), Q(3, 0), R(3, 5), S(0, 10) This trapezoid is dilated by a scale factor of 4. What is the location of point R'? A. (12, 9) B. (7, 9) C. (12, 20) D. (7, 20)

OpenStudy (anonymous):

if you were to solve to X it would be y-45 over 10=x

OpenStudy (anonymous):

what do u think? @easy! can u visualize?!

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what do u think is the answer?

OpenStudy (anonymous):

@telltoamit

OpenStudy (anonymous):

c

OpenStudy (anonymous):

what do u think is the answer? @easy!

OpenStudy (anonymous):

Correct! :)

OpenStudy (anonymous):

Which of the following relations is a function? A. (1, 4), (-4, 2), (8, 1), (-8, 2) B. (1, 4), (-4, 6), (1, 3), (-8, 2) C. (1, 0), (-4, 3), (8, 1), (-4, 5) D. (8, 1), (-4, 4), (1, 1), (8, 2)

OpenStudy (anonymous):

B

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

Which of the following best describes the equation below? y = 8x3 - 4 A. both linear and nonlinear B. neither linear nor nonlinear C. nonlinear D. linear

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

Solve the following system of equations. 8x + 5y = -70 3x + y = -21 A. x = -6, y = -5 B. x = 5, y = -22 C. x = -5, y = -6 D. x = 12, y = -57

OpenStudy (anonymous):

@telltoamit

OpenStudy (anonymous):

Use substitution of y

OpenStudy (anonymous):

@shelbyryne

OpenStudy (anonymous):

@seiga

OpenStudy (anonymous):

@tomo

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

@AriPotta

OpenStudy (anonymous):

In triangle UVW, the measure of WUV is 110° and the measure of UVW is 41°. In triangle XYZ, the measure of ZXY is 110° and the measure of YZX is 29°. Determine how triangles UVW and XYZ can be shown to be similar. A. Since WUV ZXY and UV = YZ, the triangles are similar by angle-side. B. Since WUV ZXY and UV = XZ, the triangles are similar by angle-side. C. Since WUV ZXY and UVW YZX, the triangles are similar by angle-angle. D. Since WUV ZXY and UVW XYZ, the triangles are similar by angle-angle.

OpenStudy (anonymous):

@TSwizzle

OpenStudy (anonymous):

I think its A. Btw im gunna need some medals for this!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

In triangle UVW, the measure of WUV is 110° and the measure of UVW is 41°. In triangle XYZ, the measure of ZXY is 110° and the measure of YZX is 29°. Determine how triangles UVW and XYZ can be shown to be similar. A. Since WUV ZXY and UV = YZ, the triangles are similar by angle-side. B. Since WUV ZXY and UV = XZ, the triangles are similar by angle-side. C. Since WUV ZXY and UVW YZX, the triangles are similar by angle-angle. D. Since WUV ZXY and UVW XYZ, the triangles are similar by angle-angle.

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

hay

OpenStudy (anonymous):

@ahoward79

OpenStudy (anonymous):

@Ashleyisakitty

OpenStudy (anonymous):

In triangle UVW, the measure of WUV is 110° and the measure of UVW is 41°. In triangle XYZ, the measure of ZXY is 110° and the measure of YZX is 29°. Determine how triangles UVW and XYZ can be shown to be similar. A. Since WUV ZXY and UV = YZ, the triangles are similar by angle-side. B. Since WUV ZXY and UV = XZ, the triangles are similar by angle-side. C. Since WUV ZXY and UVW YZX, the triangles are similar by angle-angle. D. Since WUV ZXY and UVW XYZ, the triangles are similar by angle-angle.

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

I think

OpenStudy (anonymous):

is it b or a

OpenStudy (anonymous):

do you know

OpenStudy (anonymous):

@EvonHowell

OpenStudy (anonymous):

@ahoward79

OpenStudy (anonymous):

@@phi

OpenStudy (anonymous):

@seiga

OpenStudy (anonymous):

If the sphere shown above has a radius of 14 units, then what is the approximate volume of the sphere? A. 1,829.33 cubic units B. 4,573.33 cubic units C. 784 cubic units D. 3,658.67 cubic units

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

@goformit100

OpenStudy (anonymous):

please help

OpenStudy (evonhowell):

I'm not currently in this math so I can't really help :(